# GroundGazer: Camera-based indoor localization of mobile robots with millimeter accuracy at low cost

### v2

@@ I Introduction
High-accuracy planar indoor localization applications, e.g. docking in warehousing and benchmarking low-accuracy indoor positioning systems with up to cm localization accuracy, require localization approaches with reliable mm positioning accuracy. Currently, such systems are very costly (generally $1000\,\mathdollar$, with costs rising quickly with increasing size of the coverage area). In this work, we introduce a novel camera-based indoor localization system for AMRs which achieves mm localization and sub-degree heading estimation accuracy with high robustness against occlusion of the position references on the ground at multitudes lower cost than existing systems of similar precision. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground.
@@ I-A Localization technologies with mm positioning accuracy
Radio frequency (RF)-based backscatter systems using frequency modulated continuous wave (FMCW) radar and reflective tags (e.g. Van-Atta arrays) that modulate the reflected radar signal can also achieve 3D mm localization accuracy using high-precision range estimates of the tags and multilateration (MLAT) [hawkeye] or triangulation for positioning [mm_siemens].
@@ III-B QR code detection and masking
Since the QR codes would disturb the successive image processing operations, we mask them out. Therefore, the average color around the detected QR code is computed and the region of the QR code is replaced with a slightly larger rectangle of this color. This masking procedure might also mask out parts of the crosshair if it lies on a QR code. The QR code masking could be improved in future work, but our experiments have shown that our system can cope with irregular missed crosshair detections. With some additional processing, we could also fully reconstruct a partly occluded crosshair.
@@ III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter after grayscale conversion and Gaussian smoothing for edge detection and mask out the previously detected laser crosshair (after extending the crosshair lines), as shown in Fig. 4(b). The binary Canny output is then processed by Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation, which consists of erosion followed by dilation. In dilation and erosion, a 2D kernel is slided over the image, where the kernel center corresponds to the pixel to process. In each kernel, erosion only keeps white (detected line) pixels if all kernel pixels are white. Dilation does the opposite and sets the color of the center pixel to white if any pixel in the kernel is white. The image after Hough transformation and opening is depicted in Fig. 4(c). For further operations, the image is inverted, as seen in Fig. 4(d). The chessboard squares are detected by first finding and simplifying the white square contours with the functions findContours [find_contours] and approxPolyDP [ramer, douglas_peuker]. The simplified contours (polygons) are then filtered for square detection. Only polygons that have four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. As the detected squares are slightly smaller due to the line broadening and opening, 10 pixels are added to the square corner coordinates. Those detected squares are given in Fig. 4(e). In the last step, the detected corners are refined to subpixel accuracy with cornerSubPix [subpix] shown in Fig. 4(f).
@@ III-E Position and heading estimation
Using this crosshair tracking after an initial global position and heading estimate with a QR code, one could theoretically avoid frequent global updates with the QR code references, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square while crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the square size, the sampling rate of the camera, or add a robust crosshair position prediction model (e.g. through odometry).

### v3

@@ I Introduction
High-accuracy planar indoor localization applications, e.g. docking in warehousing and benchmarking lower-accuracy indoor positioning systems with up to cm localization accuracy, require localization approaches with reliable mm positioning accuracy. Currently, such systems are very costly (generally $1000\,\mathdollar$, with costs rising quickly with increasing size of the coverage area). In this work, we introduce a novel camera-based indoor localization system for AMRs which achieves mm localization and sub-degree heading estimation accuracy with high robustness against occlusion of the position references on the ground at multitudes lower cost than existing systems of similar precision. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground.
@@ I-A Localization technologies with mm positioning accuracy
Radio frequency (RF)-based backscatter systems using frequency modulated continuous wave (FMCW) radar and reflective tags (e.g. Van-Atta arrays) that modulate the reflected radar signal can also achieve 3D mm localization accuracy using high-precision range estimates of the tags and multilateration (MLAT) [hawkeye] or triangulation [mm_siemens] for positioning.
@@ III-B QR code detection and masking
Since the QR codes would disturb the successive image processing operations, we mask them out. Therefore, the average color around the detected QR code is computed and the region of the QR code is replaced with a slightly larger rectangle of this color. This masking procedure might also mask out parts of the crosshair if it lies on a QR code. The QR code masking could be improved in future work, but our experiments have shown that our system can cope with irregular missed crosshair detections. By proper parametrization of the Hough transform in the crosshair detection, we can also fully reconstruct a partly occluded crosshair.
@@ III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter for edge detection after grayscale conversion and Gaussian smoothing and mask out the previously detected laser crosshair (after extending the crosshair lines), as shown in Fig. 4(b). The binary Canny output is then processed by Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation, which consists of erosion followed by dilation. In dilation and erosion, a 2D kernel is slided over the image, where the kernel center corresponds to the pixel to process. In each kernel, erosion only keeps white (detected line) pixels if all kernel pixels are white. Dilation does the opposite and sets the color of the center pixel to white if any pixel in the kernel is white. The image after Hough transformation and opening is depicted in Fig. 4(c). For further operations, the image is inverted, as seen in Fig. 4(d). The chessboard squares are detected by first finding and simplifying the white square contours with the functions findContours [find_contours] and approxPolyDP [ramer, douglas_peuker]. The simplified contours (polygons) are then filtered for square detection. Only polygons that have four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. As the detected squares are slightly smaller due to the line broadening and opening, 10 pixels are added to the square corner coordinates. Those detected squares are given in Fig. 4(e). In the last step, the detected corners are refined to subpixel accuracy with cornerSubPix [subpix] shown in Fig. 4(f).
@@ III-E Position and heading estimation
Using this crosshair tracking after an initial global position and heading estimate with a QR code, further frequent global updates with the QR code references are not necessary, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square while crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the square size, the sampling rate of the camera, or add a robust crosshair position prediction model (e.g. through odometry).